Project

General

Profile

Meeting 9-9-2015

Date: 9th Aug 2015
Time: 15:00 UK time
Attendance: Melissa, Paul, Adam, Ray Chris Heidt,
Apologies:

Agenda and Minutes

Actions from Last Meeting

  • Prepare a tag out lock out procedure to stop people turning things off that we don't want them to (He etc. - Add the above to the controls and monitoring alarms so that if the above happens we know - Pierrick. - ongoing
  • Add to John Govans' rota for daily walk round to test our setup. - Pierrick. - ongoing
  • Multiple tracks per Tracker. - Durga - ongoing
  • Calibration is in good state to go in CDB. Has not made it into last release though - C Heidt - ongoing branch being created
  • Startup and operations procedure to be added to the ISIS Tracker user document. - MU - ongoing branch being created

1. Status of hardware,

Run readiness: EO, AB, CMW, PR, MU

Since last user run the trackers have been used but not heavily.
Since last week 24/7 detector running begun - pedestal noise and LED data. -data is only available in MLCR
Data runs automatically and emails Ed if something goes wrong
Running okay but cryo 3 power cuts out intermittently - PS to be changed - ACTION
DAQ upgraded from Davids old code.
Regular calibration peds and led runs should happen throughout runs so we need to implement a structure for this with shifter instructions. - Ed to speak to Steve.
Very few errors using new DAQ code. Ed is currently working out the best analysis method. Tolerance levels needs to be determined. - ACTION
May add tolerance references to online monitoring plots.
Calibration data goodness test needs to be done - Action
Cryo2 compressor broken. Replaced by Hydrogen test system spare.

Hall Probes: MU

*Installed US
  • Running taking data since Friday

MICE hall, inc Services and requirements : CM, GB,

  • Helium fill - Diffuser rebuild 0n target for Oct 16th with a fitting date of Oct 19th and a new guy Antonio on board to help the fitting
  • Interlocks. - delayed to next available point. Must not happen too close to a user run.
  • Tracker lab tidied.

Ongoing Plan: EO, AB, CMW, PR, MU
As above

2. Software: PK, AD, CHunt, CHeidt

  • CDB going in
  • CHeidt - looking at low level data to test for sign flip. - ongoing

Geometry, Calibrations, Configurations, CDB - C. Heidt

Monte Carlo

  • In a good state waiting on Durga for multiple tracks per Tracker.

Pattern Recognition - AD,

  • No method to get multiple tracks per trigger as not available in MC.
  • Happy now with straight line track fitting. However there are still some outliers that have to be understood.
  • Helical tracks would be interesting to have but may not be in the next user run. Only 1% with more than 1 track in each tracker.

Online Displays - MU

  • Committed and in latest release. Need to add Kuno plot and info box.

Performance and Analysis - PK

Tracker Software Paper and Documentation - AD, PK DR

3. Analysis of first data

  • Update on bad channel noting/removing - Chris Heidt

4. Alignment

5. Optimising the DAQ and Controls and Monitoring - EO, AB, PR

8. Tracker Documentation

For ISIS:
We need to produce documentation for the control room and ISIS, for Steve B. We cannot necessarily get that far with operations and startup as we are not yet working with the CandM as we will but we can write something and finalise safety details.

Need to add startup and operations procedure to the document.

  • Startup Procedure
  • Operations Procedure
  • Tracker Online Emergency Procedure
  • Documentation for ISIS and eventually shifters in what to do in an emergency situation for the Tracker when running
    Documentation to Include:
    - schematics of the system
    - a list of associated checklists at power on, and who works through them.
    - a list of instructions to shifters on how to operate and monitor the system
    - emergency shutdown instructions
    - power-up instructions
    - trouble-shooting : what to watch for
    - a list of regular maintenance and safety checks

5. Other Items

  • Installation Manual - GB MU - ongoing
  • LED system - EO - External used for calib, Internal not used since Dec but will be used after gain and pedestals determined.

AoB
Next meeting will be a face to face at RAL to review progress and what needs to be done before step IV.

New Actions

  • Add 24/7 tracker data runs to grid.
  • cryo3 PS to be swapped 11/9/15
  • Add regular off MICE running led and pedestal tests to grid. - ED Paolo and Henry
  • Regular calibration peds and led runs should happen throughtout runs so we need to implement a structure for this with shifter instructions. - Ed to speak to Steve.
  • Determine tolerance levels for all calibration parameters - ED and ALL!
  • Calibration data goodness test needs to be done - Ed
  • Interlocks move. - tell Daresbury Must not happen too close to a user run. -MU
  • Need new compressor spare to replace that o=from cryo2 and new borrowed spare from H system - MU to check with Craig.

Phone Details

  • The four digit company id: 4673 (callers from outside the UK)
  • The pin number: 419357

Phone numbers:

UK (Imperial College --preferred) 0845 545 4460 English
Austria 0820 4000 1503 English
0820 4000 15 02 German
Belgium 070 35 99 45 English
070 35 99 94 Dutch
070 35 98 66 French
Canada 1 712 432 2832 English
1 712 432 2844 French
Germany 01803 001 178 English
01570 248 8303 English (mobile)
01803 001 177 German
01570 248 8101 German (mobile)
Ireland 0818 270 007 English
Italy 848 390 166 English
848 391 819 Italian
Netherlands 0870 001 901 English
0870 001 909 Dutch
Poland 0801 003 543 English
0801 003 533 Polish
South Africa 087 550 0375 English
Spain 902 885 318 English
902 881 200 Spanish
Sweden 0939 2066 400 English
0939 2066 300 Swedish
Switzerland 0848 560 190 English
0848 560 378 French
0848 560 347 German
0848 560 385 Italian
United Kingdom (PowWowNow) 0844 4 73 73 73 English
United States 1 415 363 0833 English

Bad_Channel_Finding.pdf (84.7 KB) Heidt, Christopher, 09 September 2015 14:54

Analysis.docx (7.82 MB) George, Melissa, 09 September 2015 16:20